# Cropping outperforms dropout as an augmentation strategy for training self-supervised text embeddings

### v1

@@ Abstract
Text embeddings, i.e. vector representations of entire texts, play an important role in many NLP applications, such as retrieval-augmented generation, sentiment analysis, clustering, or visualizing collections of texts for data exploration. Currently, top-performing embedding models are derived from pre-trained language models via extensive supervised fine-tuning using curated text pairs. This contrasts with computer vision, where self-supervised training based on data augmentations has demonstrated remarkable success. Here we systematically compare the two most well-known augmentation strategies for positive pair generation in contrastive learning of text embeddings. We assess embedding quality on MTEB and additional in-domain evaluations and show that cropping augmentation strongly outperforms the dropout-based approach. We find that on out-of-domain data, the quality of resulting embeddings is below the supervised SOTA models, but for in-domain data, self-supervised fine-tuning produces high-quality text embeddings after very short fine-tuning, sometimes only marginally below the supervised SOTA. Finally, we show that representation quality increases towards the last transformer layers, which undergo the largest change during fine-tuning; and that fine-tuning only those last layers is sufficient to reach similar embedding quality.
@@ 1 Introduction
Representing texts as vectors is important in natural language processing for both supervised (spam detection, sentiment analysis, semantic matching) and unsupervised (clustering, visualization, retrieval) downstream tasks. Such representations (or text embeddings) can be obtained with a wide range of methods, from simple bag-of-words representations such as TF-IDF (Jones, 1972) to transformer-based large language models (LLMs) (Zhao et al., 2023). These language models are initially trained (pre-trained) with a token-level loss, and subsequent fine-tuning with a text-level loss is needed to obtain useful text-level representations (Xu et al., 2023). We refer to models and representations fine-tuned for representing entire texts as sentence transformers and sentence or text embeddings, following Reimers & Gurevych (2019).
In recent benchmarks such as MTEB (Muennighoff et al., 2023), sentence transformers relying on extensive supervised contrastive fine-tuning on large curated datasets have typically performed best, whereas self-supervised contrastive training resulted in worse models. This is in stark contrast to computer vision, where self-supervised learning (SSL) via data augmentations has been immensely successful in producing semantically meaningful image representations (Balestriero et al., 2023). Various SSL approaches have been suggested for sentence-level fine-tuning — such as the ones employed by SimCSE (Gao et al., 2021) or DeCLUTR (Giorgi et al., 2021) — but it remains unclear how their performance compares between each other and to the state-of-the-art (SOTA) embedding models.
In this work we address these questions by systematically comparing the two most well-known data augmentation strategies used for training text embeddings, and by comparing them to the state-of-the-art supervised models. By extensively assessing representation quality on MTEB tasks and additional in-domain evaluations, we show that:
  • using text crops as positive pairs for contrastive learning performs consistently better than the dropout-based augmentation used by SimCSE, contrary to some claims in the literature (Gao et al., 2021);
  • on out-of-domain data, the quality of resulting embeddings is below the supervised SOTA models, but for in-domain data, self-supervised fine-tuning produces high-quality sentence embeddings, only marginally below supervised SOTA;
@@
  • a large part of the improvement during SSL fine-tuning is due to the generic and domain-independent sentence adaptation;
@@
Our findings are noteworthy given that dropout augmentations of SimCSE (Gao et al., 2021) represent one of the most well-cited SSL approaches in the literature on text embeddings. This augmentation strategy has also been adopted as a self-supervised training component in a recent work on transforming generative LLMs into text embedding models (BehnamGhader et al., 2024), suggesting that our findings may be relevant for ongoing work.
@@ 2 Related work
Transformer-based language models receive a sequence of text tokens as input and produce a separate latent representation for each of the tokens as output (Vaswani et al., 2017). The BERT model (Devlin et al., 2019) and its later variants such as MPNet (Song et al., 2020) include an additional classification token [CLS] to serve as a global representation of full sentences in downstream tasks. However, only a small fraction of typical BERT training is dedicated to sentence-level tasks, such that [CLS] sentence representations do not usually perform well at encoding sentence-level semantics (Thakur et al., 2021; Muennighoff et al., 2023; Jiang et al., 2022). Likewise, averaging all output tokens to obtain a sentence-level representation does not perform well either (Muennighoff et al., 2023).
To improve sentence-level representations, more sophisticated pooling strategies (Wang & Kuo, 2020) and post-processing techniques (Li et al., 2020; Su et al., 2021) have been suggested. Alternatively, a token-level model can be fine-tuned with a sentence-level objective, typically using contrastive learning (Xu et al., 2023). Here, pairs of similar texts are used as positive pairs, which are pulled together in the embedding space. Approaches differ in how similar texts are defined.
@@
In supervised contrastive learning, positive pairs are collected based on some explicit notion of similarity. Sentence-BERT (SBERT) (Reimers & Gurevych, 2019) uses a curated dataset of paired texts such as question-answer pairs from Stack Exchange; their most recent model (2021) was trained on over 1 billion of such pairs. Similarly, BGE undergoes contrastive fine-tuning using a curated dataset of text pairs (Xiao et al., 2024). For academic texts, SPECTER (Cohan et al., 2020) and SciNCL (Ostendorff et al., 2022) use citing and cited paper abstracts to form positive pairs. Sentence-T5 (ST5) (Ni et al., 2021) and Sentence-GPT (SGPT) (Muennighoff, 2022) are derived via contrastive fine-tuning of T5 (Raffel et al., 2020) and GPT (Radford et al., 2018) models.
@@
In self-supervised contrastive learning, positive pairs are generated automatically from unpaired texts, similar to self-supervised learning in computer vision that relies on data augmentations (Chen et al., 2020). The unsupervised SimCSE model from Gao et al. (2021) uses two different dropout patterns to form a positive pair of embeddings. This approach has also been used by Liu et al. (2021) and Yan et al. (2021), who additionally investigate other data augmentation techniques, such as randomly masking a part of the input text or shuffling tokens. Outputs of two distinct networks can also be used to generate positive pairs (Kim et al., 2021; Carlsson et al., 2021). Further, one can use adjacent chunks of a text as positive pairs; this was applied to train RNN (Logeswaran & Lee, 2018), GPT (Neelakantan et al., 2022), and BERT models (Giorgi et al., 2021; Izacard et al., 2022). Recently, synthetic generation of positive pairs has been explored leveraging generative LLMs (Zhang et al., 2023), but this approach comes with the additional computational cost of text generation.
@@
In past benchmarks of sentence transformers, models trained in a supervised way have been shown to outperform the ones trained with self-supervision (Thakur et al., 2021; Muennighoff et al., 2023). For example, SBERT’s latest all-mpnet-base-v2 achieved top results among all models of BERT-base size. The BGE-base embedding model (Xiao et al., 2024) and its further modifications are currently best performing on MTEB among models of BERT-base size. Larger models, such as BGE-large or commercial embedding models like text-embedding-3-large from OpenAI and embed-english-v3 from Cohere, outperform BERT-base-sized models in some evaluation tasks, but in others are on par with smaller models.
@@ 3.1 Augmentations and loss function
We set out to investigate how effectively sentence representations from a language model can be improved through self-supervised fine-tuning alone, and how different augmentation choices influence this improvement. For that, we leveraged a contrastive learning approach analogous to SimCLR (Chen et al., 2020) as our training strategy and compared various augmentation techniques for generating positive pairs, such as text crops (Logeswaran & Lee, 2018; Giorgi et al., 2021; Neelakantan et al., 2022), dropout-based augmentation (Gao et al., 2021), and variations of those (see Section A.1). Importantly, all augmentation approaches in our comparison were constructed in a self-supervised manner, without relying on external notions of similarity or annotated data.
@@
Sentence embeddings, i.e. vector representations of entire texts, play an important role in many NLP applications, such as retrieval augmented generation, sentiment analysis, clustering, or visualizing collections of texts for data exploration. Currently, top-performing embedding models are derived from pre-trained language models via extensive supervised fine-tuning. $\leftrightarrow\,\,$ This contrasts with computer vision, where self-supervised training has demonstrated remarkable success. Here we study whether good sentence representations can be obtained through self-supervised fine-tuning alone, assessing embedding quality on MTEB and via additional in-domain evaluations.
@@
where $\text{sim}(\bm{a},\bm{p})=\bm{a}^{\top}\bm{p}/\big{(}\|\bm{a}\|\cdot\|\bm{p}\|\big{)}$ is the cosine similarity between $\bm{a}$ and $\bm{p}$, the vector representations of texts $a$ and $p$. We set the temperature to $\tau=0.05$ and the batch size to $b=64$, the largest possible batch size given our GPU memory resources. We trained the network using the Adam optimizer (Kingma & Ba, 2014) with learning rate $\eta=2\cdot 10^{-5}$, with linear warm-up and linear decay. See Section A.1 for details on hyperparameter choices.
@@ 3.2.1 Setup
We fine-tuned a base transformer (i.e., a model primarily pre-trained with a token-level loss) using the SSL training setup outlined in the previous section and compared two augmentation strategies: cropping and dropouts. We chose MPNet (Song et al., 2020; mpnet-base) as base model, following SBERT (Reimers & Gurevych, 2019; model all-mpnet-base-v2 on HuggingFace). This model has bert-base architecture with 110 M parameters and uses 768 embedding dimensions. We used mean pooling over all tokens to obtain a single 768-dimensional output vector for each input text, but also compared other pooling strategies (Supplementary Figure S1).
For fine-tuning the base model we used the ICLR dataset (González-Márquez & Kobak, 2024), which consists of 24,347 scientific abstracts of all papers submitted to the ICLR conference within the years 2017–2024. This recently assembled dataset is well-suited for our purposes because it is not part of the MTEB benchmark. Fine-tuning on this dataset therefore enables a fair model comparison and avoids evaluation data leakage, which can potentially lead to inflated performance estimates.
For simplicity, we did not use all MTEB tasks for evaluation, but focused on a subset of tasks from four different modalities: clustering, reranking, retrieval, and STS. Clustering tasks assess the $K$-means clustering results in the embedding space; retrieval and reranking tasks assess the quality of the nearest neighbors in the embedding space; STS tasks measure how well the embedding represents not only small but also large ground-truth pairwise distances (see Section A.2 for details). These evaluation modalities quantify different aspects of sentence representation and cover the wide range of MTEB tasks.
@@ 3.2.2 Results
Table 1: MTEB tasks. Row blocks correspond to clustering, reranking, retrieval, and STS tasks. All values in percent, higher is better. Models in columns 3–4 were fine-tuned on the ICLR dataset.
@@
ArxivClusteringP2P 27.8 35.4 33.3 38.3 48.1 48.6 48.7
BiorxivClusteringP2P 23.2 30.1 31.1 32.4 39.3 39.7 39.4
MedrxivClusteringP2P 22.5 28.0 29.3 30.8 35.6 32.6 33.2
RedditClusteringP2P 37.4 44.7 49.5 55.9 56.6 64.7 62.7
StackExchangeCl…P2P 26.3 28.8 30.2 31.3 34.3 35.0 35.2
SciDocsRR 56.1 69.5 64.6 73.6 88.7 87.6 87.5
MindSmallReranking 27.5 29.3 28.4 30.2 31.0 32.1 31.2
SCIDOCS 1.4 7.9 6.5 13.0 23.8 22.6 21.7
ArguAna 22.2 41.4 41.9 50.6 46.5 64.5 63.8
@@
STS16 50.6 77.7 66.2 76.0 80.0 86.5 85.5
STSBenchmark 52.0 78.6 67.9 71.7 83.4 87.5 86.4
Block average 33.2 46.8 42.8 48.7 55.2 58.7 58.2
@@
Out of the box, MPNet resulted in poor representations with a block average across modalities of 33.2% (Table 1, column 1). After fine-tuning on the ICLR dataset for a single epoch, the quality of the embeddings markedly improved. Out of the two augmentation strategies, cropping worked much better than dropout: cropping-based fine-tuning outperformed dropout-based fine-tuning in all tasks (on average 48.7% vs 42.8%, Table 1, columns 3–4). Cropping-based fine-tuning yielded an improvement in block average score of 14.8 percentage points, but there were large differences between modalities: STS had the largest improvement (21.2 p.p.), followed by retrieval (19.4 p.p.), reranking (10.1 p.p.), and clustering (8.6 p.p.).
@@
The off-the-shelf unsupervised SimCSE model (which is based on dropout fine-tuning) performed similarly to our dropout-based fine-tuned model on all tasks except for the STS, where it was substantially better (Table 1, column 2 and 3), suggesting that the good performance of SimCSE in STS benchmarks may be due to some other fine-tuning choices beyond the dropout augmentation. SimCSE yielded worse results than our cropping-based fine-tuning in the other three modalities, despite having being fine-tuned on two orders of magnitude more data (1 M Wikipedia sentences for SimCSE vs. 24 k samples in our experiments), demonstrating that, at least for most tasks, the choice of augmentation has a greater impact on performance than the amount of training data.
@@
Cropping-based fine-tuned MPNet was 6.5 percentage points below SBERT, which achieved 55.2% performance as block average. Fine-tuned MPNet performed closest to SBERT in reranking tasks (3.3 p.p. below SBERT) and furthest in STS tasks (9.6 p.p. below SBERT). This demonstrates that cropping-based fine-tuning produced a sentence-level model that showed substantial generalization despite very limited amount of fine-tuning ($\sim$24 k training samples, ICLR dataset only). Note also that some of the evaluation datasets, e.g. SCIDOCS, arXiv, and StackExchange, formed part of the training set of SBERT, possibly biasing SBERT performance estimates upwards.
@@
Additionally, we compared the standard transformer models to two state-of-the-art models: BGE-base (109 M parameters) and BGE-large (335 M parameters) (Xiao et al., 2024) (see Table S1 for the list of all used models). BGE-base is among the best-performing models in public MTEB leaderboard within models of its size. BGE-large is the larger version of BGE-base, and is among the top models overall, on par with text-embedding-3-large from OpenAI. In the MTEB tasks we evaluated, both BGE models performed slightly better than SBERT (3 p.p. higher in block average) and substantially better than our cropping-based fine-tuned MPNet (10 p.p. difference). This suggests that additional training modifications, such as pre-training specifically targeted at sentence embeddings, instruction-based supervised fine-tuning, or hard negative samples, can bring embedding quality further.
Figure 1: Fine-tuning on ICLR dataset.

### v2

# Cropping outperforms dropout as an augmentation strategy for self-supervised training of text embeddings

@@ Abstract
Text embeddings, i.e. vector representations of entire texts, play an important role in many NLP applications, such as retrieval-augmented generation, clustering, or visualizing collections of texts for data exploration. Currently, top-performing embedding models are derived from pre-trained language models via supervised contrastive fine-tuning. This fine-tuning strategy relies on an external notion of similarity and annotated data for generation of positive pairs. Here we study self-supervised fine-tuning and systematically compare the two most well-known augmentation strategies used for fine-tuning text embeddings models. We assess embedding quality on MTEB and additional in-domain evaluations and show that cropping augmentation strongly outperforms the dropout-based approach. We find that on out-of-domain data, the quality of resulting embeddings is substantially below the supervised state-of-the-art models, but for in-domain data, self-supervised fine-tuning can produce high-quality text embeddings after very short fine-tuning. Finally, we show that representation quality increases towards the last transformer layers, which undergo the largest change during fine-tuning; and that fine-tuning only those last layers is sufficient to reach similar embedding quality.
@@ 1 Introduction
Representing texts as vectors is important in natural language processing for both supervised (e.g. spam detection, sentiment analysis, semantic matching) and unsupervised (e.g. clustering, visualization, retrieval) downstream tasks. With the advent of retrieval-augmented generation (RAG), it has become increasingly important to produce high-quality text representations that enable retrieval of relevant texts that can be added to the prompt to improve performance of generative large language models (LLMs) (Izacard et al., 2022b; Ram et al., 2023). Such representations (or text embeddings) can be obtained with a wide range of methods, from simple bag-of-words representations such as TF-IDF (Jones, 1972) to transformer-based large language models (see Zhao et al., 2023, for a survey). Such language models are initially trained (pre-trained) with a token-level loss, and subsequent fine-tuning with a text-level loss is needed to obtain useful text-level representations (Xu et al., 2023). We refer to models and representations fine-tuned for representing entire texts as sentence transformers and sentence or text embeddings, following Reimers and Gurevych (2019).
In recent benchmarks such as MTEB (Muennighoff et al., 2023), sentence transformers relying on extensive supervised contrastive fine-tuning on large curated datasets have typically performed best, whereas models solely trained with self-supervised contrastive learning scored worse results. This is in stark contrast to computer vision, where self-supervised learning (SSL) via data augmentations has been immensely successful in producing semantically meaningful image representations (Balestriero et al., 2023). Various SSL approaches have been suggested for sentence-level fine-tuning, including the dropout-based method employed by SimCSE (Gao et al., 2021) and the cropping-based method employed by DeCLUTR (Giorgi et al., 2021). However, it remains unclear how their performance compares between each other and to the state-of-the-art (SOTA) embedding models, especially when evaluated on texts from some specific limited domain.
In this work we address these questions by systematically comparing the two most well-known data-augmentation strategies used for training text embeddings, and by comparing them to the state-of-the-art supervised models. By extensively assessing representation quality on MTEB tasks and additional in-domain evaluations, we show that:
  • using text crops as positive pairs for contrastive learning performs consistently better than the dropout-based augmentation used by SimCSE, contrary to the claims in Gao et al. (2021);
  • on out-of-domain data, the quality of resulting embeddings is substantially below the supervised SOTA models, but for in-domain data, self-supervised fine-tuning produces high-quality sentence embeddings, comparable to those by the supervised SOTA models;
@@
  • a large part of the improvement during SSL fine-tuning is due to the generic and domain-independent adaptation that can be called sentence adaptation;
@@
Our findings are noteworthy given that dropout augmentations of SimCSE (Gao et al., 2021) represent one of the most well-known and well-cited SSL approaches in the literature on text embeddings [^1] At the moment of publication, Gao et al. (2021) has 4 900 citations in Google Scholar, second only after Reimers and Gurevych (2019) with 22 300 citations, in the literature on text/sentence embeddings.. Dropout augmentations are still being used to fine-tune modern models (e.g. BehnamGhader et al., 2024) and have also been adopted in other fields, e.g. for representation learning of single-cell RNA-sequencing data (Yang et al., 2022).
We therefore believe that, even though modern text-embedding models are usually relying on supervised fine-tuning, systematically comparing different augmentation approaches for self-supervised fine-tuning can provide valuable insights into which training strategies yield the best text representations.
@@ 2 Related work
2.1 Text embeddings from generative and non-generative language models
Early text embeddings were obtained with bag-of-words approaches, such as TF-IDF (Jones, 1972) and BM25 (Robertson et al., 1995; Kamphuis et al., 2020). Transformer-based (Vaswani et al., 2017) language models led to a shift toward dense embeddings, which capture richer semantic information and context compared to the sparse bag-of-words representations. Initially, non-generative encoder-only (i.e. BERT-like) models were used for generating text embeddings since they use bidirectional attention, which is more appropriate for embedding tasks. However, recent work has increasingly adopted generative decoder-only (i.e. GPT-like) models for embedding tasks to leverage powerful recent models with more extensive pre-training.
Encoder-only models
@@ 2.1 Text embeddings from generative and non-generative language models
These models receive a sequence of text tokens as input and produce a separate latent representation for each of the tokens as output, without any restrictions on the attention patterns, resulting in bidirectional self-attention. The BERT model (Devlin et al., 2019) and its later variants such as RoBERTa (Liu et al., 2019) and MPNet (Song et al., 2020) include an additional classification token [CLS] to serve as a global representation of the full text in downstream tasks. However, only a small fraction of typical BERT training is dedicated to sentence-level tasks, such that [CLS] representations do not usually perform well at encoding sentence-level semantics (Thakur et al., 2021; Muennighoff et al., 2023). Simply averaging the embeddings across all tokens to obtain a sentence-level representation does not perform well either (Muennighoff et al., 2023). Therefore, more sophisticated pooling (Wang and Kuo, 2020) and post-processing strategies (Li et al., 2020; Su et al., 2021) have been suggested to improve BERT-derived sentence representations. Alternatively, a model pre-trained on the token level can be fine-tuned with a sentence-level objective to improve the sentence-level representations (see below). Usually fine-tuning is performed on the average pooling of all tokens or the [CLS] token embedding.
Decoder-only models
Decoder-only models are used for text generation and employ masked (as opposed to bidirectional) attention, where each token can only attend to preceding tokens. Different strategies have been proposed to adapt generative models for text-embedding tasks. One approach is to append an instruction before the text and then use either the representation of the last [EOS] token (Wang et al., 2024; Li et al., 2025) or the average representation across all tokens (excluding the instruction) (BehnamGhader et al., 2024) as the text embedding. This strategy was originally introduced in encoder-only models (Su et al., 2023), but has increasingly been adopted for generative models. A generative model (with or without additional instructions) can also be fine-tuned with a sentence-level objective (see below). Prior to such fine-tuning, some works remove the causal attention masking, switching the model to bidirectional attention (BehnamGhader et al., 2024; Lee et al., 2025). Additionally, some fine-tuned models use a particular way of pooling such as position-weighted mean pooling (Muennighoff, 2022) or latent attention pooling (Lee et al., 2025).
Note that some tasks that are often used for benchmarking text embeddings can be handled by a generative model without any explicit text embeddings. For example, a classification task can be presented to a generative LLM with a prompt asking to predict the text’s label from a given set of labels (Chen et al., 2025). This does not constitute a text embedding.
2.2 Contrastive fine-tuning of text embeddings
Whatever the exact model architecture and pre-training are, contrastive learning is commonly used for fine-tuning of text representations. Contrastive learning pulls together pairs of similar texts (positive pairs) and pushes apart pairs of non-similar texts (negative pairs) in the embedding space (Chen et al., 2020). Contrastive learning approaches differ in how similar texts are defined.
Supervised contrastive learning
@@ 2.2 Contrastive fine-tuning of text embeddings
Here positive pairs are collected based on some explicit notion of similarity. Sentence-BERT (SBERT) (Reimers and Gurevych, 2019) uses a curated dataset of paired texts such as question-answer pairs from Stack Exchange; their most recent model all-mpnet-base-v2 (2021) was trained on over one billion of such pairs. Similarly, BGE (Xiao et al., 2024), E5 (Wang et al., 2022), and GTE (Li et al., 2023) all undergo contrastive fine-tuning using large datasets of text pairs curated from different sources, such as internet webpages (title-body pairs), scientific papers (title-abstract pairs), or community forums like StackExchange (question-answer pairs). For academic texts, SPECTER (Cohan et al., 2020) and SciNCL (Ostendorff et al., 2022) use citing and cited papers’ abstracts to form positive pairs. Sentence-T5 (ST5) (Ni et al., 2022) and Sentence-GPT (SGPT) (Muennighoff, 2022) are derived via contrastive fine-tuning of T5 (Raffel et al., 2020) and GPT (Radford et al., 2018) models on curated datasets of text pairs.
Self-supervised contrastive learning
@@
Here positive pairs are generated automatically from unpaired texts, similar to self-supervised learning in computer vision that relies on data augmentations (Chen et al., 2020). SimCSE (Gao et al., 2021) uses two different dropout patterns to form a positive pair of embeddings. This approach has also been used by Liu et al. (2021) and Yan et al. (2021) who additionally investigate other data-augmentation techniques such as randomly masking parts of the input text. The same dropout approach was also adopted by LLM2Vec (BehnamGhader et al., 2024) for fine-tune generative LLMs (after removing attention masking, see above). Alternatively, outputs of two distinct networks can be used to generate positive pairs (Kim et al., 2021; Carlsson et al., 2021). Further, one can use adjacent chunks of a text as positive pairs; this was applied to train RNN (Logeswaran and Lee, 2018), GPT (Neelakantan et al., 2022), and BERT models (DeCLUTR, Giorgi et al., 2021; Contriever, Izacard et al., 2022a; and CoCondenser, Gao and Callan, 2022). Recently, synthetic generation of positive pairs has been explored leveraging generative LLMs (Zhang et al., 2023; Wang et al., 2024), but this approach comes with the additional computational cost of text generation.
2.3 Benchmarks
@@
In benchmarks of sentence transformers, models trained in a supervised way outperform the ones trained only with self-supervision (Thakur et al., 2021; Muennighoff et al., 2023). Among the models of BERT-base size, the BGE-base model (Xiao et al., 2024) with its further modifications and the GTE-base model (Li et al., 2023) are currently among the best performers, overtaking the SBERT’s latest all-mpnet-base-v2 model. Larger models, such as BGE-large or commercial embedding models like text-embedding-3-large from OpenAI and embed-english-v3 from Cohere, outperform BERT-base-sized models, in particular on some of the evaluation tasks.
@@ 3.1 Augmentations and loss function
We set out to investigate how effectively sentence representations from a language model can be improved through self-supervised fine-tuning alone, and how different augmentation choices influence this improvement. For that, we leveraged a contrastive learning approach analogous to SimCLR (Chen et al., 2020) as our training strategy and compared various augmentation techniques for generating positive pairs, such as text crops (Logeswaran and Lee, 2018; Giorgi et al., 2021; Neelakantan et al., 2022; Izacard et al., 2022a; Gao and Callan, 2022), dropout-based augmentation (Gao et al., 2021), and variations of those (see Section A.1). Importantly, all augmentation approaches in our comparison were constructed in a self-supervised manner, without relying on external notions of similarity or annotated data.
@@
Text embeddings, i.e. vector representations of entire texts, play an important role in many NLP applications, such as retrieval-augmented generation, clustering, or visualizing collections of texts for data exploration. Currently, top-performing embedding models are derived from pre-trained language models via supervised contrastive fine-tuning. $\leftrightarrow\,\,$ This fine-tuning strategy relies on an external notion of similarity and annotated data for generation of positive pairs. Here we study self-supervised fine-tuning and systematically compare the two most well-known augmentation strategies used for fine-tuning text embeddings models.
@@
where $\text{sim}(\bm{a},\bm{p})=\bm{a}^{\top}\bm{p}/\big(\|\bm{a}\|\cdot\|\bm{p}\|\big)$ is the cosine similarity between $\bm{a}$ and $\bm{p}$, the vector representations of texts $a$ and $p$. We set the temperature to $\tau=0.05$ and the batch size to $b=64$, the largest possible batch size given our GPU memory resources. We trained the network using the Adam optimizer (Kingma and Ba, 2015) with learning rate $\eta=2\cdot 10^{-5}$, with linear warm-up and linear decay. See Section A.1 for details on hyperparameter choices.
@@ 3.2.1 Setup
We fine-tuned a base transformer (i.e. a model primarily pre-trained with a token-level loss) using the SSL training setup outlined in the previous section and compared two augmentation strategies: cropping and dropouts. We chose MPNet (Song et al., 2020; mpnet-base) as base model, following SBERT (Reimers and Gurevych, 2019; model all-mpnet-base-v2 on HuggingFace). This model has bert-base architecture with 110 M parameters and uses 768 embedding dimensions. We used mean pooling over all tokens to obtain a single 768-dimensional output vector for each input text, but also compared other pooling strategies (Section A.1, Supplementary Figure S1). Furthermore, we also tested two other base models (BERT and RoBERTa; Devlin et al., 2019; Liu et al., 2019; Supplementary Table S1).
For fine-tuning the base model we used the ICLR dataset (González-Márquez and Kobak, 2024), which consists of 24,347 scientific abstracts of all papers submitted to the ICLR conference within the years 2017–2024. This recently assembled dataset is well-suited for our purposes because it is not part of the MTEB benchmark. Fine-tuning on this dataset therefore enables a fair model comparison and avoids evaluation data leakage, which can potentially lead to inflated performance estimates.
For simplicity, we did not use all MTEB tasks for evaluation, but focused on a subset of English tasks from five different modalities: clustering, reranking, retrieval, STS, and classification. Clustering tasks assess the $K$-means clustering results in the embedding space; retrieval and reranking tasks assess the quality of the nearest neighbors in the embedding space; STS tasks measure how well the embedding represents not only small but also large ground-truth pairwise distances; classification tasks use a logistic regression classifier to assess linear separability of ground-truth classes (see Section A.2 for details). These evaluation modalities quantify different aspects of sentence representation and cover the wide range of MTEB tasks. We used SentenceTransformers library that normalizes all text-embedding vectors to have unit norm.
@@ 3.2.2 Results
Table 1: MTEB tasks. Row blocks correspond to clustering, reranking, retrieval, STS, and classification tasks. All values in percent, higher is better. Models in columns 3–4 were fine-tuned on the ICLR dataset.
@@
ArxivClusteringP2P 27.8 35.4 33.3 38.3 48.1 48.7 48.6
BiorxivClusteringP2P 23.2 30.1 31.1 32.4 39.3 39.4 39.7
MedrxivClusteringP2P 22.5 28.0 29.3 30.8 35.6 33.2 32.6
RedditClusteringP2P 37.4 44.7 49.5 55.9 56.6 62.7 64.7
StackExchangeCl…P2P 26.3 28.8 30.2 31.3 34.3 35.2 35.0
SciDocsRR 56.1 69.5 64.6 73.6 88.7 87.5 87.6
MindSmallReranking 27.5 29.3 28.4 30.2 31.0 31.2 32.1
SCIDOCS 1.4 7.9 6.5 13.0 23.8 21.7 22.6
ArguAna 22.2 41.4 41.9 50.6 46.5 63.8 64.5
@@
STS16 50.6 77.7 66.2 76.0 80.0 85.5 86.5
STSBenchmark 52.0 78.6 67.9 71.7 83.4 86.4 87.5
AmazonPolarityClass. 66.5 72.0 63.3 62.5 67.1 93.4 92.4
Banking77Class. 57.4 74.4 63.3 67.7 81.7 87.0 87.8
ImdbClass. 61.8 66.5 66.2 64.9 71.2 90.8 92.8
MassiveIntentClass. 55.9 60.5 53.0 53.9 69.8 72.6 74.3
MassiveScenarioClass. 60.7 66.7 62.8 65.9 75.7 76.5 77.4
MTOPDomainClass. 75.9 82.8 82.6 85.1 91.9 93.2 94.0
TweetSent…Class. 52.8 52.2 51.6 49.7 55.0 59.4 59.9
Block average 38.9 51.0 46.9 51.8 58.8 62.9 63.5
@@
Out of the box, MPNet resulted in poor representations with a block average across modalities of 38.9% (Table 1, column 1). After fine-tuning on the ICLR dataset for a single epoch, the quality of the embeddings markedly improved. Out of the two augmentation strategies, cropping worked much better than dropout: cropping-based fine-tuning outperformed dropout-based fine-tuning in 16/19 tasks (on average 51.8% vs 46.9%, Table 1, columns 3–4), and in the other 3 the difference was small. Cropping-based fine-tuning yielded an improvement in block average score of 12.9 percentage points, but there were large differences between modalities: STS had the largest improvement (21.4 p.p.), followed by retrieval (20.0 p.p.), clustering (10.3 p.p.), reranking (10.1 p.p.), and classification (2.7 p.p).
We repeated the same experiment using two further base models (BERT and RoBERTa) and observed the same results as with MPNet, namely that fine-tuning using cropping augmentation yielded better representations than fine-tuning using dropout (Supplementary Table S1). Representations obtained with cropping-based fine-tuning were better than those obtained via dropout in 16/19 tasks for BERT and in 15/19 for RoBERTa. Compared to the out-of-the-box models, after cropping-based fine-tuning the representations improved by 6.8 percentage points for BERT and by 10.7 p.p. for RoBERTa. The exact ranking of MTEB modalities by improvement size varied between models, but retrieval always showed one of the most pronounced improvements, while in classification tasks there was near-zero change when using BERT and RoBERTa.
@@
The off-the-shelf unsupervised SimCSE model (which is based on dropout fine-tuning) performed similarly to our dropout-based fine-tuned model on all tasks except for the STS and some of the classification tasks, where it was substantially better than all fine-tuned base models (Table 1, columns 2 and 3; Table S1, columns 5 and 8), suggesting that the good performance of SimCSE in those benchmarks may be due to some other fine-tuning choices beyond the dropout augmentation. SimCSE yielded worse results than our cropping-based fine-tuning in the other three modalities, despite having being fine-tuned on two orders of magnitude more data (1 M Wikipedia sentences for SimCSE vs. 24 k samples in our experiments), demonstrating that, at least for most tasks, the choice of augmentation has a greater impact on performance than the amount of training data.
@@
Cropping-based fine-tuned MPNet was 7.0 percentage points below SBERT, which achieved 58.8% performance as block average. Fine-tuned MPNet performed closest to SBERT in retrieval tasks (3.3 p.p. below SBERT) and furthest in STS tasks (9.6 p.p. below SBERT). This demonstrates that cropping-based fine-tuning produced a sentence-level model that showed substantial generalization despite very limited amount of fine-tuning ($\sim$24 k training samples, ICLR dataset only). Note also that some of the evaluation datasets, e.g. SCIDOCS, arXiv, and StackExchange, formed part of the training set of SBERT, possibly biasing SBERT performance estimates upwards.
@@
Additionally, we compared the standard transformer models to two state-of-the-art models: BGE-base (109 M parameters) and BGE-large (335 M parameters) (Xiao et al., 2024) (see Table S2 for the list of all used models). BGE-base is among the best-performing models in public MTEB leaderboard within models of its size. BGE-large is the larger version of BGE-base, and is among the top models overall, on par with text-embedding-3-large from OpenAI. In the MTEB tasks we evaluated, both BGE models performed better than SBERT, especially on classification tasks (4.4 p.p. higher in block average) and substantially better than our cropping-based fine-tuned MPNet (11.4 p.p. difference). This suggests that additional training modifications, such as pre-training specifically targeted at sentence embeddings, instruction-based supervised fine-tuning, or hard negative samples, can bring embedding quality further.